### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in modern medicine, yet their application in Traditional Chinese Medicine (TCM) remains severely limited by the absence of standardized benchmarks and the scarcity of high-quality training data. To address these challenges, we introduce TCM-Eval, the first dynamic and extensible benchmark for TCM, meticulously curated from national medical licensing examinations and validated by TCM experts. Furthermore, we construct a large-scale training corpus and propose Self-Iterative Chain-of-Thought Enhancement (SI-CoTE) to autonomously enrich question-answer pairs with validated reasoning chains through rejection sampling, establishing a virtuous cycle of data and model co-evolution. Using this enriched training data, we develop ZhiMingTang (ZMT), a state-of-the-art LLM specifically designed for TCM, which significantly exceeds the passing threshold for human practitioners. To encourage future research and development, we release a public leaderboard, fostering community engagement and continuous improvement.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has catalyzed a paradigm shift across numerous domains Zhao et al. (2023); Gao et al. (2025); Wang et al. (2024), with healthcare emerging as a particularly promising frontier Thirunavukarasu et al. (2023); Zhou et al. (2023). These models have demonstrated remarkable capabilities in tasks such as clinical decision support Li et al. (2025); Garza et al. (2025), medical text summarization Van Veen et al. (2024); Aali et al. (2025), and patient-facing conversational AI Mukherjee et al. (2024), significantly accelerating innovation in the medical field.
However, despite these advancements, the focus of current medical LLMs and open-source datasets has been overwhelmingly skewed towards modern medicine Liu et al. (2024b); Wu et al. (2024); Chen et al. (2024); Zhang et al. (2025). This has created a significant gap in the domain of Traditional Chinese Medicine (TCM). The development of powerful LLMs for TCM is severely hampered by two fundamental challenges: the absence of a unified, authoritative platform for model evaluation Chen et al. (2025), and the scarcity of high-quality, large-scale training data Qiu et al. (2024); Zhang et al. (2024). This lack of foundational resources restricts progress and prevents a systematic understanding of current models’ capabilities within the TCM domain.
@@
To address the challenge of scarce high-quality training data, we have compiled a comprehensive corpus from 18 authoritative TCM textbooks and thousands of mock examinations, which we used to create a domain-specific dataset consisting of over 384,807 question-answer pairs. To leverage this data and enhance reasoning abilities, we introduce Self-Iterative Chain-of-Thought Enhancement (SI-CoTE), designed to autonomously augment simple QA pairs with high-quality, step-by-step reasoning processes. The SI-CoTE process operates iteratively: the model generates potential reasoning chains for a subset of the data, and through Rejection Sampling Liu et al. (2023), we retain only those chains that lead to the correct answer. This newly generated high-quality CoT data is then used to fine-tune the model, yielding a more capable version. This enhanced model, in turn, processes the next subset of data, creating a virtuous cycle where both the training data quality and the model’s reasoning abilities evolve synergistically. Using this efficient self-improvement strategy, we developed ZhiMingTang (ZMT) by fine-tuning the powerful Deepseek-R1 DeepSeek-AI et al. (2025) foundation model.
We conducted extensive experiments on TCM-Eval using both open-source and closed-source models. ZMT achieved an impressive average score of 96.32, significantly outperforming all 34 competing models. It also far exceeds the 60-point passing score required for human practitioners in the official examination. This demonstrates that ZMT has attained expert-level knowledge and reasoning abilities in TCM, matching or even surpassing the capabilities of qualified human professionals. Additionally, we observed a significant data leakage issue in current models, where performance on newly added questions was notably lower compared to older questions. In contrast, ZMT showed no such decline, maintaining consistent performance across both old and new questions.
@@
  • We build a high-quality QA dataset from mock exams and authoritative TCM textbooks. Using the SI-CoTE method, we expanded it into QA pairs with reasoning chains. This enriched dataset is used to train ZMT, establishing a new SOTA in the TCM domain.
@@ 2 Related Work
Datasets Dynamic TCMLE- Textbook Real Mock Expert- Tasks Scale
@@
TCM-Bench Yue et al. (2024) MCQ & MTQ 5,473
TCM-3CEval Huang et al. (2025) MCQ 450
MTCMB Kong et al. (2025) QA & MCQ & FIB 7,100
TCMEval-SDT Wang et al. (2025) SQA 300
TCM-Ladder Xie et al. (2025) MCQ & FIB & QA & VQA 52K
OphthBench Zhou et al. (2025) MCQ & QA 591
TCM-Eval(Ours) MCQ 6,099
Table 1: Overview of Datasets in the TCM Domain. TCMLE-Specifc refers to datasets specifically designed for the TCM Licensing Examination. MCQ instands for Multiple-choice Question, MTQ instands for Matching-type Questions, FIB instands for Fill-in-the-blank Question, MQA stands for Multi-modal Question Answering, and SQA stands for Structured Question Answering.
@@ 2.1 Large Language Models for Traditional Chinese Medicine
One type of research integrates knowledge from external TCM knowledge bases into open-source large language models to assist the models in generating responses. TCM-KLLaMA Zhuang et al. (2025) extracts knowledge from knowledge graphs and injects it into the input text of the model; BenTsao Wang et al. (2023), OpenTCM He et al. (2025) utilizes knowledge graph based retrieval enhanced generation technology to explicitly utilize knowledge from the knowledge graph bases during inference.
@@
One type of research applies simple question-answer pairs for Supervised Fine-Tuning (SFT) Ding et al. (2023) on open-source large language models, enabling the models to acquire TCM domain knowledge. TCMLLM Haoyu et al. (2024), TCM-FTP Zhou et al. (2024), ShenNong TCM-LLM Zhu et al. (2023), BianCang Sibo et al. (2024), TCMChat Dai et al. (2024) are fine-tuned on instructions based on TCM knowledge and Chinese question answering; BianQue Chen et al. (2023) , Qibo Jia et al. (2025) and CMLM-ZhongJing Kang et al. (2025) build multi-turn dialogue datasets based on doctor-patient roles and inquiry scenarios to enhance their consultation capabilities.
@@
The other type of research adopts chain-of-thought (CoT) Wei et al. (2023) data for SFT on open-source large language models, allowing the models to not only grasp the basic knowledge of the TCM domain but also possess the reasoning capabilities of TCM domain experts. Qibo Jia et al. (2025) and Lingdan Hua et al. (2024) construct Chain-of-Thought for TCM consultation prompts, endowing the models with preliminary thinking logic and reasoning capabilities; JingFang Yang et al. (2025b) improves the model’s ability in comprehensive clinical consultation and precise syndrome differentiation through the design of Multi-Agent Collaborative Chain-of-Thought Mechanism (MACCTM) and Dual-Stage Recovery Scheme (DSRS).
However, these methods only apply static CoT data to prompts for model reasoning, failing to form dynamic iteration of CoT data.In contrast, our ZMT autonomously enriches question-answer pairs with validated reasoning chains through rejection sampling, establishing a virtuous cycle of data and model co-evolution.
@@ 2.2 Datasets for Traditional Chinese Medicine
In recent years, evaluation benchmarks specifically designed for the TCM domain have emerged. TCMBench Yue et al. (2024) targets practitioner exams across 16 knowledge areas, while TCM-3CEval Huang et al. (2025) and MTCMB Kong et al. (2025) broaden the scope to include literature understanding, diagnostic reasoning, and prescription recommendation. Benchmarks such as TCMEval-SDT Wang et al. (2025) and TCM-Ladder Xie et al. (2025) emphasize structured and multimodal reasoning, with the latter incorporating hierarchical difficulty and text–image inputs. Comprehensive frameworks like TCMBench Yue et al. (2024) cover both theory and clinical decision-making, and cross-domain efforts such as OphthBench Zhou et al. (2025) demonstrate adaptability to specialized fields. However, most existing benchmarks are static and narrow in coverage limiting their long-term applicability. In contrast, the TCM-Eval dataset introduces a dynamic, expert-validated benchmark based on NMLE-TCM, with a systematically constructed training corpus derived from 18 authoritative textbooks and mock exams.
@@ 3 Data Collection
Our proposed dataset can be divided into two parts: TCM-Eval, a benchmark for TCM developed with expert oversight and inter-rater consistency to ensure accuracy(Section 3.1); and TCM-Corpus, a training corpus constructed via an expert–LLM collaborative pipeline enhanced by our Self-Iterative Chain-of-Thought Enhancement (SI-CoTE) framework, which was used to train our model (Section 3.2). We apply rigorous quality control protocols to both parts (Section 3.3), and report their distributions, domain coverage, and key statistics (Section 3.4).
@@ 3.1 TCM-Eval
The National Qualification Examination for TCM Practitioners, recognized as the authoritative professional certification in China, integrates the full scope of TCM across 18 official textbooks, making it the field’s de facto gold standard. TCM-Eval is designed to rigorously evaluate the reasoning, memorization, and application abilities of models in the TCM domain by leveraging this gold standard. Unlike the TCM-Corpus, which emphasizes knowledge acquisition and exposure to diverse question formats, TCM-Eval exclusively prioritizes authenticity, reliability, and dynamic renewal to ensure fair and representative benchmarking.
@@ 3.2 TCM-Corpus
The domain of TCM encompasses not only a broad spectrum of knowledge (including both classical TCM theories and modern medical concepts ) but also constitutes a self-contained system of dialectical reasoning, which introduces unique challenges for model training. Drawing inspiration from the authentic learning process of human students, we design a data collection pipeline that systematically extracts knowledge-oriented question–answer pairs from textbooks and acquires mock examination items through automated procedures, thereby emulating the dual processes of study and practice in TCM education. Furthermore, we propose a self-iterative framework to generate high-quality Chain-of-Thought (CoT) data.
@@
To training our reasoning model — ZhiMingTang, we proposed Self-Iterative Chain-of-Thought Enhancement (SI-CoTE) framework. This framework is built upon the Deepseek-R1 Guo et al. (2025) base model, denoted as $\mathcal{M}_{0}$. At its core, the SI-CoTE methodology employs a phased, iterative process to efficiently convert the original Question-Answer (QA) pair dataset, $\mathcal{D}_{\text{QA}}$, into an enhanced dataset containing high-quality CoT data, thereby driving the evolution of the model’s capabilities.
@@ 3.3 Quality Control
For TCM-Corpus, a model-in-the-loop validation procedure was introduced Liu et al. (2020); Cheng et al. (2025); Liu et al. (2025). Items consistently answered correctly by the model were retained as high-confidence samples, while items answered incorrectly were flagged for stricter review, including both automated heuristics and human inspection. This hybrid validation process yielded a final curated dataset.
@@ 3.4 Data Statistics
We conducted a detailed statistical analysis of the training and test sets for TCM-Eval, as shown in Table 3 and Table 2. The training set covers 16 knowledge domains in TCM, constructed from 1,601 sets of practice questions and 18 authoritative official textbooks, comprising a total of 384,807 instruction-tuning samples and 159 million tokens. These data enable the model to acquire foundational knowledge and key examination topics in TCM during the post-training phase.
@@ 4.1 Setup
Following previous work Chen et al. (2025); Cheng et al. (2025), we conducted extensive experiments on a wide range of open-source and API-based models.
Specifically, the open-sourced models encompass various series with different parameter scales. For general-purpose large language models, our selection includes Qwen3 series (8B, 14B, 32B) Yang et al. (2025a), Llama-3 series (8B, 70B) Grattafiori et al. (2024), Deepseek series (V3.1, R1) Liu et al. (2024a); Guo et al. (2025), Baichuan-M2 (32B) Dou et al. (2025), and Mistral-8B [^1] https://mistral.ai/news/ministraux, GPT-oss series (20B, 120B) OpenAI (2025), the DS-Qwen series (7B, 14B, 32B), and the DS-Llama series (8B, 70B). To account for domain-specific applications, we also incorporated models known for their performance in the medical field, namely the MedGemma series (4B, 27B) Sellergren et al. (2025) and the ShiZhenGPT series (7B, 32B) Chen et al. (2025).
For the API-based models, our evaluation includes services from various providers. These include GLM-4 series (GLM-4.5, GLM-4.5-Air) Team et al. (2025a), Kimi-K2-Instruct Team et al. (2025b), Baichuan4222https://platform.baichuan-ai.com/, Ernie-x1-turbo-32k , Spark-4.0-Ultra [^3] https://xinghuo.xfyun.cn/, MiniMax-M1 MiniMax et al. (2025), LongCat-Flash-Chat Team (2025), and GPT-4 series [^4] https://chatgpt.com/ (GPT-4o, GPT-4.1).
@@
To ensure the reproducibility and fairness of our results, we standardized the inference and fine-tuning procedures for all models. For inference, we employed the efficient vLLM framework Kwon et al. (2023) as the unified engine for all open-sourced models. For non-reasoning or deterministic generation tasks, we set the decoding temperature to 0 with both top-p and top-k configured to 0.1, whereas for tasks requiring reasoning, the temperature was set to 0.6. In the Supervised Fine-Tuning (SFT) phase, we customized the SFTTrainer from the Hugging Face Transformers library Wolf et al. (2020) and adopted the parameter-efficient Low-Rank Adaptation (LoRA) Hu et al. (2022). All experiments were conducted on a single node equipped with eight H20 GPUs, each with 141 GB of VRAM, and the entire training phase consumed approximately 7,000 GPU-hours.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable capabilities in modern medicine, yet their application in Traditional Chinese Medicine (TCM) remains severely limited by the absence of standardized benchmarks and the scarcity of high-quality training data. To address these challenges, we introduce TCM-Eval, the first dynamic and extensible benchmark for TCM, meticulously curated from national medical licensing examinations and validated by TCM experts. Furthermore, we construct a large-scale training corpus and propose Self-Iterative Chain-of-Thought Enhancement (SI-CoTE) to autonomously enrich question-answer pairs with validated reasoning chains through rejection sampling, establishing a virtuous cycle of data and model co-evolution. Using this enriched training data, we develop ZhiMingTang-M1 (ZMT-M1), a state-of-the-art LLM specifically designed for TCM, which significantly exceeds the passing threshold for human practitioners. To encourage future research and development, we release a public leaderboard, fostering community engagement and continuous improvement.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has catalyzed a paradigm shift across numerous domains zhao2023survey; gao2025survey; survey_data_synthesis, with healthcare emerging as a particularly promising frontier thirunavukarasu2023large; zhou2023survey. These models have demonstrated remarkable capabilities in tasks such as clinical decision support li2025questionsclinicalrecommendationslarge; garza2025retrievalaugmentedframeworkllmbasedclinical, medical text summarization van2024adapted; aali2025dataset, and patient-facing conversational AI mukherjee2024polaris, significantly accelerating innovation in the medical field.
However, despite these advancements, the focus of current medical LLMs and open-source datasets has been overwhelmingly skewed towards modern medicine liu2024medbenchcomprehensivestandardizedreliable; wu2024medjourney; chen2024clinicalbench; zhang2025llmeval. This has created a significant gap in the domain of Traditional Chinese Medicine (TCM). The development of powerful LLMs for TCM is severely hampered by two fundamental challenges: the absence of a unified, authoritative platform for model evaluation chen2025shizhengptmultimodalllmstraditional, and the scarcity of high-quality, large-scale training data qiu2024buildingmultilinguallanguagemodel; zhang2024qibolargelanguagemodel. This lack of foundational resources restricts progress and prevents a systematic understanding of current models’ capabilities within the TCM domain.
@@
To address the challenge of scarce high-quality training data, we have compiled a comprehensive corpus from 18 authoritative TCM textbooks and thousands of mock examinations, which we used to create a domain-specific dataset consisting of over 384,807 question-answer pairs. To leverage this data and enhance reasoning abilities, we introduce Self-Iterative Chain-of-Thought Enhancement (SI-CoTE), designed to autonomously augment simple QA pairs with high-quality, step-by-step reasoning processes. The SI-CoTE process operates iteratively: the model generates potential reasoning chains for a subset of the data, and through Rejection Sampling liu2023statistical, we retain only those chains that lead to the correct answer. This newly generated high-quality CoT data is then used to fine-tune the model, yielding a more capable version. This enhanced model, in turn, processes the next subset of data, creating a virtuous cycle where both the training data quality and the model’s reasoning abilities evolve synergistically. Using this efficient self-improvement strategy, we developed ZhiMingTang-M1 (ZMT-M1) by fine-tuning the powerful Deepseek-R1 deepseekai2025deepseekr1incentivizingreasoningcapability foundation model.
We conducted extensive experiments on TCM-Eval using both open-source and closed-source models. ZMT-M1 achieved an impressive average score of 96.32, significantly outperforming all 34 competing models. It also far exceeds the 60-point passing score required for human practitioners in the official examination. This demonstrates that ZMT-M1 has attained expert-level knowledge and reasoning abilities in TCM, matching or even surpassing the capabilities of qualified human professionals. Additionally, we observed a significant data leakage issue in current models, where performance on newly added questions was notably lower compared to older questions. In contrast, ZMT-M1 showed no such decline, maintaining consistent performance across both old and new questions.
@@
  • We build a high-quality QA dataset from mock exams and authoritative TCM textbooks. Using the SI-CoTE method, we expanded it into QA pairs with reasoning chains. This enriched dataset is used to train ZMT-M1, establishing a new SOTA in the TCM domain.
@@ 2 Related Work
Datasets Dynamic TCMLE- Textbook Real Mock Expert- Scale
@@
TCM-Bench yueTCMBenchComprehensiveBenchmark2024 5,473
TCM-3CEval huangTCM3CEvalTriaxialBenchmark2025 450
MTCMB kongMTCMBMultiTaskBenchmark2025 7,100
TCMEval-SDT wangTCMEvalSDTBenchmarkDataset2025 300
TCM-Ladder xieTCMLadderBenchmarkMultimodal2025 52K
OphthBench zhouOphthBenchComprehensiveBenchmark2025 591
TCM-Eval(Ours) 6,099
Table 1: Overview of Datasets in the TCM Domain. TCMLE-Specific refers to datasets specifically designed for the TCM Licensing Examination.
@@ 2.1 Large Language Models for Traditional Chinese Medicine
One type of research integrates knowledge from external Traditional Chinese Medicine (TCM) knowledge bases into open-source large language models to assist the models in generating responses. TCM-KLLaMA ZHUANG2025109887 extracts knowledge from knowledge graphs and injects it into the input text of the model; BenTsao wang2023huatuo, OpenTCM he2025opentcmgraphragempoweredllmbasedtraditional utilizes knowledge graph based retrieval enhanced generation technology to explicitly utilize knowledge from the knowledge graph bases during inference.
@@
One type of research applies simple question-answer pairs for Supervised Fine-Tuning (SFT) Ding2023ParameterefficientFO on open-source large language models, enabling the models to acquire TCM domain knowledge. TCMLLM haoyu2024tcmllm, TCM-FTP zhou2024tcmftpfinetuninglargelanguage, ShenNong TCM-LLM zhu2023ChatMed, BianCang Wei2024BianCang, TCMChat DAI2024107530 are fine-tuned on instructions based on TCM knowledge and Chinese question answering; BianQue chen2023bianque , Qibo JIA2025127672 and CMLM-ZhongJing Kang2025 build multi-turn dialogue datasets based on doctor-patient roles and inquiry scenarios to enhance their consultation capabilities.
@@
The other type of research adopts chain-of-thought (CoT) wei2023chainofthoughtpromptingelicitsreasoning data for SFT on open-source large language models, allowing the models to not only grasp the basic knowledge of the TCM domain but also possess the reasoning capabilities of TCM domain experts. Qibo JIA2025127672 and Lingdan hua2024lingdan construct Chain-of-Thought for TCM consultation prompts, endowing the models with preliminary thinking logic and reasoning capabilities; JingFang yang2025jingfangexpertlevellargelanguage improves the model’s ability in comprehensive clinical consultation and precise syndrome differentiation through the design of Multi-Agent Collaborative Chain-of-Thought Mechanism (MACCTM) and Dual-Stage Recovery Scheme (DSRS).
However, these methods only apply static CoT data to prompts for model reasoning, failing to form dynamic iteration of CoT data.In contrast, our ZMT-M1 autonomously enriches question-answer pairs with validated reasoning chains through rejection sampling, establishing a virtuous cycle of data and model co-evolution.
@@ 2.2 Datasets for Traditional Chinese Medicine
In recent years, evaluation benchmarks specifically designed for the TCM domain have emerged. TCMBench yueTCMBenchComprehensiveBenchmark2024 targets practitioner exams across 16 knowledge areas, while TCM-3CEval huangTCM3CEvalTriaxialBenchmark2025 and MTCMB kongMTCMBMultiTaskBenchmark2025 broaden the scope to include literature understanding, diagnostic reasoning, and prescription recommendation. Benchmarks such as TCMEval-SDT wangTCMEvalSDTBenchmarkDataset2025 and TCM-Ladder xieTCMLadderBenchmarkMultimodal2025 emphasize structured and multimodal reasoning, with the latter incorporating hierarchical difficulty and text–image inputs. Comprehensive frameworks like TCMBench yueTCMBenchComprehensiveBenchmark2024 cover both theory and clinical decision-making, and cross-domain efforts such as OphthBench zhouOphthBenchComprehensiveBenchmark2025 demonstrate adaptability to specialized fields. However, most existing benchmarks are static and narrow in coverage limiting their long-term applicability. In contrast, the TCM-Eval dataset introduces a dynamic, expert-validated benchmark based on NMLE-TCM, with a systematically constructed training corpus derived from 18 authoritative textbooks and mock exams.
@@ 3 Data Collection
Our proposed dataset can be divided into two parts: TCM-Eval, a benchmark for Traditional Chinese Medicine (TCM) developed with expert oversight and inter-rater consistency to ensure accuracy(Section 3.1); and TCM-Corpus, a training corpus constructed via an expert–LLM collaborative pipeline enhanced by our Self-Iterative Chain-of-Thought Enhancement (SI-CoTE) framework, which was used to train our model (Section 3.2). We apply rigorous quality control protocols to both parts (Section 3.3), and report their distributions, domain coverage, and key statistics (Section 3.4).
@@ 3.1 TCM-Eval
The National Qualification Examination for Traditional Chinese Medicine (TCM) Practitioners, recognized as the authoritative professional certification in China, integrates the full scope of TCM across 18 official textbooks, making it the field’s de facto gold standard. TCM-Eval is designed to rigorously evaluate the reasoning, memorization, and application abilities of models in the TCM domain by leveraging this gold standard. Unlike the TCM-Corpus, which emphasizes knowledge acquisition and exposure to diverse question formats, TCM-Eval exclusively prioritizes authenticity, reliability, and dynamic renewal to ensure fair and representative benchmarking.
@@ 3.2 TCM-Corpus
The domain of Traditional Chinese Medicine (TCM) encompasses not only a broad spectrum of knowledge ( including both classical TCM theories and modern medical concepts ) but also constitutes a self-contained system of dialectical reasoning, which introduces unique challenges for model training. Drawing inspiration from the authentic learning process of human students, we design a data collection pipeline that systematically extracts knowledge-oriented question–answer pairs from textbooks and acquires mock examination items through automated procedures, thereby emulating the dual processes of study and practice in TCM education. Furthermore, we propose a self-iterative framework to generate high-quality Chain-of-Thought (CoT) data.
@@
To training our reasoning model — ZhiMingTang, we proposed Self-Iterative Chain-of-Thought Enhancement (SI-CoTE) framework. This framework is built upon the Deepseek-R1 guo2025deepseek base model, denoted as $\mathcal{M}_{0}$. At its core, the SI-CoTE methodology employs a phased, iterative process to efficiently convert the original Question-Answer (QA) pair dataset, $\mathcal{D}_{\text{QA}}$, into an enhanced dataset containing high-quality CoT data, thereby driving the evolution of the model’s capabilities.
@@ 3.3 Quality Control
For TCM-Corpus, a model-in-the-loop validation procedure was introduced liu-etal-2020-towards-conversational; cheng-etal-2025-toolspectrum; liu2025repodebugrepositorylevelmultitaskmultilanguage. Items consistently answered correctly by the model were retained as high-confidence samples, while items answered incorrectly were flagged for stricter review, including both automated heuristics and human inspection. This hybrid validation process yielded a final curated dataset.
@@ 3.4 Data Statistics
We conducted a detailed statistical analysis of the training and test sets for TCM-Eval, as shown in Table 3 and Table 2. The training set covers 16 knowledge domains in traditional Chinese medicine, constructed from 1,601 sets of practice questions and 18 authoritative official textbooks, comprising a total of 384,807 instruction-tuning samples and 159 million tokens. These data enable the model to acquire foundational knowledge and key examination topics in TCM during the post-training phase.
@@ 4.1 Setup
Following previous work chen2025shizhengptmultimodalllmstraditional; cheng-etal-2025-toolspectrum, we conducted extensive experiments on a wide range of open-source and API-based models.
Specifically, the open-sourced models encompass various series with different parameter scales. For general-purpose large language models, our selection includes Qwen3 series (8B, 14B, 32B) yang2025qwen3, Llama-3 series (8B, 70B) grattafiori2024llama, Deepseek series (V3.1, R1) liu2024deepseek; guo2025deepseek, Baichuan-M2 (32B) dou2025baichuan, and Mistral-8B [^1] https://mistral.ai/news/ministraux, GPT-oss series (20B, 120B) openai2025gptoss120bgptoss20bmodel, the DS-Qwen series (7B, 14B, 32B), and the DS-Llama series (8B, 70B). To account for domain-specific applications, we also incorporated models known for their performance in the medical field, namely the MedGemma series (4B, 27B) sellergren2025medgemmatechnicalreport and the ShiZhenGPT series (7B, 32B) chen2025shizhengptmultimodalllmstraditional.
For the API-based models, our evaluation includes services from various providers. These include GLM-4 series (GLM-4.5, GLM-4.5-Air) 5team2025glm45agenticreasoningcoding, Kimi-K2-Instruct kimiteam2025kimik2openagentic, Baichuan4222https://platform.baichuan-ai.com/, Ernie-x1-turbo-32k , Spark-4.0-Ultra [^3] https://xinghuo.xfyun.cn/, MiniMax-M1 minimax2025minimaxm1scalingtesttimecompute, LongCat-Flash-Chat meituan2025longcatflashtechnicalreport, and GPT-4 series [^4] https://chatgpt.com/ (GPT-4o, GPT-4.1).
Benchmarks
As introduced in Section 3.1, we constructed TCM-Eval to comprehensively evaluate the knowledge and reasoning capabilities of LLMs in the field of TCM. We use this as our primary benchmark, testing the accuracy of each model on different question sets.
@@
To ensure the reproducibility and fairness of our results, we standardized the inference and fine-tuning procedures for all models. For inference, we employed the efficient vLLM framework kwon2023efficientmemorymanagementlarge as the unified engine for all open-sourced models. For non-reasoning or deterministic generation tasks, we set the decoding temperature to 0 with both top-p and top-k configured to 0.1, whereas for tasks requiring reasoning, the temperature was set to 0.6. In the Supervised Fine-Tuning (SFT) phase, we customized the SFTTrainer from the Hugging Face Transformers library wolf-etal-2020-transformers and adopted the parameter-efficient Low-Rank Adaptation (LoRA) hu2022lora. All experiments were conducted on a single node equipped with eight H20 GPUs, each with 141 GB of VRAM, and the entire training phase consumed approximately 7,000 GPU-hours.